A physical pendulum consists of a meter stick that is pivoted at a small hole drilled through the stick a distance x from the 50 cm mark. The period of oscillation is observed to be 2.9 s. Find the distance x.
m
Answer #1

answer) the period is given by

T=2piI/

I=moment of inertia=ML2/12 +Mx2

=Mgx

now using the above eqn we have

T2=22pi2I/

2.92=4pi2(ML2/12 +Mx2)/Mgx

2.92gx=4pi2(1/12+x2)

82.418x=3.2865+39.4384x2

39.4384x2-82.418x+3.2865=0

its in the form of quadratic eqn

A=39.4384,B=-82.418,C=3.2865

x=-bb2-4ac/2a

x=-(-82.418)-82.4182-4*39.4348*3.2865/2*29.4384

x=2.05 m or x=0.041 m

we need value less than 1 m , so first value cancel out

so the answer is 0.041 m or 0.0407m